Two different forecasting techniques (F1 and F2) were used to forecast demand for cases of bottled water. Actual demand and the two sets of forecasts are as follows:

PREDICTED DEMAND
Period Demand F1 F2
1 68 67 64
2 75 70 60
3 70 75 70
4 74 71 72
5 69 71 73
6 72 65 76
7 80 71 75
8 78 77 85


a. Compute MAD for each set of forecasts. Given your results, which forecast appears to be more accurate? (Round your answers to 2 decimal place.)

MAD F1
MAD F2


(Click to select)  F2  F1  None  appears to be more accurate.

b. Compute the MSE for each set of forecasts. Given your results, which forecast appears to be more accurate? (Round your answers to 2 decimal places.)

MSE F1
MSE F2


(Click to select)  F2  F1  None  appears to be more accurate.



d. Compute MAPE for each data set. Which forecast appears to be more accurate? (Round your intermediate calculations to 2 decimal places and and final answers to 2 decimal places.)

MAPE F1
MAPE F2

(Click to select)  F2  F1  None  appears to be more accurate.
Answer #1

FORECAST ERROR

F1

F2

BETTER?

MAD

4.13

5.13

F1

MSE

45.13

15.13

F2

MAPE

5.5

7

F1

EXPLANATION

FORECAST 1


PERIOD

ACTUAL DEMAND

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

ABS DEV / DEMAND

1

68

67

1

1

0.014706

2

75

70

5

5

0.066667

3

70

75

-5

5

0.071429

4

74

71

3

3

0.040541

5

69

71

-2

2

0.028986

6

72

65

7

7

0.097222

7

80

71

9

9

0.1125

8

78

77

1

1

0.012821


MAD = SIGMA(SUM OF ABSOLUTE DEVIATIONS) / N

MSE = SIGMA(SUM OF DEVIATIONS) ^ 2 / N

MAPE = (SIGMA(ABS DEV / DEMAND) / N) * 100

Where N represents the number of observations.

We know that N = 8

Sum of absolute deviations = 1 + 5 + 5 + 3 + 2 + 7 + 9 + 1 = 33

MAD = 33 / 8 = 4.13

Sum of deviations = 1 + 5 + -5 + 3 + -2 + 7 + 9 + 1 = 19

MSE = 19 ^2 / 8 = 45.13

Sum of abs dev / demand = 0.01 + 0.07 + 0.07 + 0.04 + 0.03 + 0.1 + 0.11 + 0.01 = 0.44

MAPE = (0.44 / 8) * 100 = 5.50


FORECAST 2


PERIOD

ACTUAL DEMAND

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

ABS DEV / DEMAND

1

68

64

4

4

0.058824

2

75

60

15

15

0.2

3

70

70

0

0

0

4

74

72

2

2

0.027027

5

69

73

-4

4

0.057971

6

72

76

-4

4

0.055556

7

80

75

5

5

0.0625

8

78

85

-7

7

0.089744



We know that N = 8

Sum of absolute deviations = 4 + 15 + 0 + 2 + 4 + 4 + 5 + 7 = 41

MAD = 41 / 8 = 5.13

Sum of deviations = 4 + 15 + 0 + 2 + -4 + -4 + 5 + -7 = 11

MSE = 11 ^2 / 8 = 15.13

Sum of abs dev/ demand = 0.06 + 0.2 + 0 + 0.03 + 0.06 + 0.06 + 0.06 + 0.09 = 0.56

MAPE = (0.56 / 8) * 100 = 7.00

**Please rate with a thumbs up. In case you have a question or query, please reach out in the comments section. Thank You**